OpenStudy (nurali):

Question 1 The elements in the circuit have the following values: R=5kΩ and rm=15kΩ Assuming that iN1=1mA and iN2=0A, what is the value of vOUT in Volts? Assuming that iN1=0A and iN2=1mA, what is the value of vOUT in Volts? Assuming that iN1=1mA and iN2=1mA, what is the value of vOUT in Volts? Assuming that Input 2 is left as an open circuit, what is the Thevenin equivalent resistance (in kOhms) seen from Input 1?

OpenStudy (anonymous):

try the formula,, for part 1 : iA = 2*i1*R/(2*R+rm) for part 2 : iA = i2*R/(2*R+rm) Part 3 is : iA(part 1)+iA(part 2)

OpenStudy (anonymous):

part 4 is R+R OR R+rm

OpenStudy (anonymous):

and what wer the answers u got

OpenStudy (anonymous):

R=3kΩ and rm=12kΩ the formulas obtained 0.00033 0.000166 0.000496 6k all wrong tell me what I'm doing wrong?